# Gökdeniz Gülmez Machine Learning Research Stuttgart, Germany

### v1

@@ 7.3 Ablation Study: Pairing Methods
  • Fisher LDA: Compute $\mathbf{w}=\mathbf{S}_{w}^{-1}(\boldsymbol{\mu}_{h}-\boldsymbol{\mu}_{n})$ where $\mathbf{S}_{w}=\frac{1}{2}(\mathbf{C}_{h}+\mathbf{C}_{n})$ is pooled within-class covariance
  • Logistic Probe: Train $p(y|\mathbf{h})=\sigma(\mathbf{w}^{\top}\mathbf{h}+b)$ on harmful/harmless labels, extract $\mathbf{w}$
@@
Results (Table 1): Across all models, SVD-pairing achieved refusal reduction $\Delta\rho=-0.87\pm 0.03$ (mean $\pm$ std over 5 models), statistically indistinguishable from Fisher LDA ($\Delta\rho=-0.89\pm 0.04$, paired t-test $p=0.12$). However, SVD required 60% less wall-clock time due to avoiding the $\mathcal{O}(d^{3})$ covariance inversion. Logistic probes underperformed ($\Delta\rho=-0.72\pm 0.05$), likely due to overfitting on the 1024-sample training set.
Table 1: Refusal rate reduction ($\Delta\rho$, lower is better) and computation time for different direction extraction methods. All methods use $k=2$ directions and identical post-extraction modification ($\alpha=0.3$, $\lambda=0.1$).
Method $\Delta\rho$ Time (s) MMLU $\Delta$
SVD-Pairing (Ours) $0.87\pm 0.03$ $42\pm 5$ $1.2\pm 0.4$
Fisher LDA $0.89\pm 0.04$ $05\pm 12 1.1$$1\pm 0.5$
Logistic Probe $0.72\pm 0.05$ $67\pm 8$ $2.8\pm 0.7$
Mean Difference $0.64\pm 0.06$ $8\pm 1$ $0.9\pm 0.3$
Conclusion: While Fisher LDA is theoretically optimal for linear discriminant analysis, the SVD-based pairing achieves comparable practical performance with significantly lower computational cost. The stochastic averaging over multiple shuffles (3–5 runs) stabilizes the extracted subspace, mitigating the apparent arbitrariness of individual pairings.
7.4 Ablation Study: Exact vs Regularized Projection
@@ Results:
  • Exact orthogonalization: Achieved strong refusal reduction ($\Delta\rho=-0.94$) but caused severe MMLU degradation ($8.3\%$ on Qwen2.5–7B, $6.1\%$ on Llama3–8B). Generated text exhibited repetition and incoherence.
  • Regularized Gabliteration: Achieved $\Delta\rho=-0.87$ with minimal MMLU degradation ($1.2\%$), maintaining fluent generation.

### v2

@@ 7.3 Ablation Study: Pairing Methods
Results (Table 1): Across all eight evaluated models, Gabliteration (Ours) exhibited lower average KL divergence, indicating reduced distributional drift and better preservation of the original model behavior while lovering the average refusal rates as well. Abliteration showed significantly higher variance across models, suggesting reduced robustness and sensitivity to model scale and architecture.
Table 1: Per-model refusal rate and KL divergence for Gabliteration and Abliteration. All models are evaluated on identical datasets (mlabonne/harmful_behaviors vs. mlabonne/harmless_alpaca), using the same refusal keywords, 100 trials, 100 harmful evaluation prompts, and 400 samples for direction extraction. Abliteration has been applied using the heretic package by Philipp Emanuel Weidmann [heretic]. Lower is better for both metrics.
Model Gabliteration (Ours) Abliteration (Mean Diff)
Refusal (%) KL Refusal (%) KL
Qwen/Qwen3–4B-Instruct-2507 4 0.1522 21 0.4300
Qwen/Qwen3–4B-Thinking-2507 3 0.0014 4 0.0500
google/gemma-3–1b-it 3 0.2922 3 0.4300
Qwen/Qwen3–0.6B 3 0.0127 6 0.0957
meta-llama/Llama-3.2–1B-Instruct 7 0.0038 8 0.0230
tencent/HY-MT1.5–1.8B 4 0.0029 15 0.0573
ibm-granite/granite-3.3–2b-instruct 6 0.0030 4 0.2100
Nanbeige/Nanbeige4–3B-Thinking-2511 8 0.0074 11 0.0024
Conclusion: Gabliteration provides a robust and scalable alternative to mean-difference-based Abliteration for refusal suppression. Across diverse model families and sizes, it achieves significantly stronger refusal reduction while inducing less distributional distortion, as measured by KL divergence. These results indicate that structured, multi-directional weight modification is critical for stable behavioral control, whereas single-direction mean-difference approaches are prone to high variance and unintended behavioral side effects.
7.4 Ablation Study: Exact vs. Regularized Projection
@@ Results:
  • Exact projection: While this approach achieved aggressive refusal suppression, it frequently introduced instability in generation, including repetition, loss of coherence, and brittle responses. These effects indicate excessive removal of representational subspaces critical for general language modeling.
  • Regularized Gabliteration: Maintained strong refusal suppression while preserving fluent and coherent generation. The regularization term effectively constrained the update magnitude, preventing catastrophic interference and yielding more stable behavior across prompts.